# iFANnpp: Nuclear Power Plant Digital Twin for Robots and Autonomous Intelligence

### v2

@@ IV-E Robot Navigation Utilities
iFANnpp includes functionality for visualizing the path that a robot takes. This feature is useful for robot debugging purposes, allowing a practitioner to visually compare the robot’s expected path against the path it took in practice (See Figure 4). Researchers can also identify potential patterns in the path that the robot is taking, improving the interpretability of its learned algorithm and directing the practitioner toward potential algorithmic or mechanical modeling modification to improve the robot’s path-planning and execution performance.
The path visualization feature is implemented in Unreal Engine as a component that can be added to a robot actor. A boolean variable can be selected through the editor to enable or disable the visualization feature. While enabled, the robot path is rendered onto the map. The robot path is implemented using a series of debug lines generated via calls to Unreal Engine ”DrawDebugLine”. Successive points along the robot path are linearly interpolated to produce a smooth visual path that trails behind the robot. Both the thickness and color of the path can be customized through the editor.
Additionally, a boolean variable is exposed through the Unreal Engine editor to allow for recording the robot location. The current robot location is recorded for every tick within the simulator and is logged to a CSV file along with the current timestamp in milliseconds. This logging feature can plot and visualize the robot path offline, independent of the iFANnpp environment, which can aid in post-hoc algorithm analysis. This feature can also be used to collect robot path demonstration trajectories to train robot agents, e.g., using behavioral cloning. Both human and machine policies can be used to produce demonstrations, as the robot can either be manually or pragmatically controlled.
To enhance the interpretability of the robot path in the first-person perspective, a compass feature is developed. This feature can also aid practitioners in reviewing and debugging first-person robot navigation task experimentation. The compass features can be enabled or disabled interactively by pressing a keyboard key while the digital twin simulation is running. This displays an overhead icon that tracks and highlights the target relative to the robot, indicating the relative angle of the target (see Figure 4). This can inform a practitioner on how to potentially adjust their algorithm to improve performance on a navigation task.
@@ IV-F Real-Time Top-Down Perspective
A top-down camera class was created within Unreal Engine, inheriting from the camera actor class. This camera is positioned at a location in which the entire map can be captured. It becomes active through a single key input. The top-down feature allows for a dynamic change in the camera view, switching from the default first-person perspective to a top-down perspective. This change in viewpoint is crucial since it provides a comprehensive overview of the environment, as shown in Figure 5. The top-down images are captured and saved automatically at predefined intervals, ensuring that critical moments and movements are recorded.
This feature offers several benefits. First, the top-down feature significantly simplifies tracing the robot’s path. This perspective allows for a clear visualization of the route taken by the robot, aiding in understanding the robot’s navigation and movement patterns. Second, this feature enables the precise identification and localization of multiple robots operating simultaneously within the environment. The functionality serves its purpose in scenarios where coordination and interaction between several robots are critical. Third, the feature allows for comprehensive process monitoring of the entire nuclear power plant. The ability to view the environment from above allows the tracking of the interactions between various elements in complex operations. This is particularly useful for ensuring the efficiency and effectiveness of the tested system or algorithm. Lastly, the feature not only includes the automatic saving of top-down images but also captures a series of images throughout the operation at predefined intervals. This establishes a continuous record of the robot’s activities, which can be used for post-operation analysis, troubleshooting, and optimizing future performance.
@@ IV-G Thermal Vision
A thermal vision feature was implemented using material functions from Unreal Engine. Thermal values are linearly interpolated between $-100$ to $100$ Celsius, visually represented by interpolating from blue to red. Actors without temperature properties are visually distinguished by being rendered in green, effectively removing them from the blue and red spectrum. This color exclusion process eliminates confusion with Unreal Engine actors that are naturally red or blue. As shown in Figure 6, this method provides a clear thermal view of the circulating water system, allowing users to interpret the temperature distribution easily. With real-time mirroring, this feature dynamically reflects the constantly changing environment and displays the entire map in thermal view.
This feature has many potential applications. First, the thermal view provides an intuitive and immediate understanding of temperature distributions across the plant. As mentioned, robots can make quicker decisions without the need to physically walk up to individual sensors. Second, this feature can enable comprehensive plant monitoring, ensuring that all critical components are consistently observed. Third, thermal images can be synthetically generated to address the lack of open-source visual thermal data in nuclear power plants. Moreover, the generated thermal data can uncover patterns, trends, and correlations that are difficult to observe in tabular datasets. While tabular data often requires thorough analysis to identify subtle trends, this feature captures complex and minor variations in the environment, enabling the perception of multiple nuanced patterns simultaneously.
@@ V Use Cases
In the following sections, a representative set of realistic scenarios is provided, utilizing iFANnpp for future research projects. These scenarios demonstrate multiple research directions that could take advantage of multiple features provided by the iFANnpp platform.
@@ V-A Example Scenario: Swarm of Robots
One valuable avenue for robotics research in nuclear power plants (as well as industrial systems in general) is swarm robotics. Robotics swarms allow for the orchestration of a relatively large number of simplistic robots capable of coordinating together to perform complex behavior. Robotic swarm deployment can be used to form flexible functional configurations, create mobile assembly lines, and perform coordinated maintenance and repair tasks.
To demonstrate the viability of simulating a robotic swarm, a swarm of twenty TurtleBots was placed within the turbine hall of the iFANnpp environment. The robots were placed adjacent to one another in a closely clustered configuration. By setting net mode within the Unreal Editor to ”Play as Listen Server” and extending the number of players to 2, multiple turtle-bots were controlled using the mouse and ”wasd” keys. Control was transferred from one robot to another.
@@
Figure 7 displays the swarm robot simulation. The left-most and right-most images display first-person camera views from two individual robots within the swarm. Each robot was individually controlled from these views. The middle image is an overhead display of the twenty-bot swarm. As with other simulation tasks, iFANnpp allows for iterating on both the mechanical and algorithmic components of swarm robot systems.
V-B Example Scenario: Robot Remote Control Via Digital Twin (teleopration)
The implementation of VR technology within the digital twin, especially in critical infrastructure like nuclear power plants, provides significant value. In the scenario provided, a teleoperation setup is demonstrated in the turbine hall, utilizing VR to allow an operator to control a robot remotely. See figure 8 for a depiction of the scenario.
@@ V-B Example Scenario: Robot Remote Control Via Digital Twin (teleopration)
GPWR is used to simulate an emergency in the circulating water system. This system is crucial for cooling, and its failure could result in failure of the overall power plant. In this case, the user navigates the robot via VR to locate and close a valve to prevent the system from over-flooding.
This application demonstrates how VR can increase safety, enabling operators to address emergencies without being physically present in dangerous areas. Furthermore, VR could help improve operator response times through enhanced situational awareness and reduced cognitive load due to immersive visualization.
There are multiple research directions that can utilize the integrating of VR technology with nuclear power plants. Human-robot collaboration and disaster preparedness simulations are relevant topics not only in the nuclear industry but also in various industrial systems. Additional applications include remote maintenance task execution and training program development.
@@ V-C Example Scenario: Robot Navigation
Research has focused extensively on enhancing the use of robots in industrial systems, mainly focusing on their movement and decision-making capabilities. One critical feature is their ability to navigate effectively, which involves selecting optimal paths while avoiding hazardous areas, such as those with high radiation levels in nuclear power plants. While robot teleportation allows for manual human control, autonomous robot navigation can be valuable when a human is not present or able to operate a robot remotely (e.g., if it is difficult to establish wireless communication). In addition, autonomous navigation capabilities can remove the burden on human operators for mundane mobile tasks such as fetching a tool that the operator needs. Therefore, research investigating the design and improvement of autonomous robot navigation algorithms is of interest.
The viability of training and deploying a RL algorithm for the task of robot path navigation in the nuclear power plant was demonstrated. The purpose of the demonstration was to determine whether a robot policy can be trained to navigate from a starting location to a target location.
The environment for this robot task was set within the reactor building. A single radiation source was placed between the robot and the target location, depicted as a colored mesh sphere. The starting point and the end point were fixed as shown in Figure 9. For each episode, the radiation source was randomly turned on or off (red indicated that the source was ”on”, blue indicated it was ”off”).
Figure 9: Reinforcement Learning Robot Navigation Scenario
DQN [34] was used in order to train a model to navigate from one part of the map to another. An ImageNet-pretrained [35] ResNet50 [36] model was used as the backbone and a 10 layer Multi-Layer Perceptron (MLP) with Leaky Rectified Linear Unit (Leaky ReLU) activation function was used to generate the approximate Q function prediction. Each hidden layer of the MLP was size 4096.
The Python-Unreal bridge was used to train the model. The client queried the iFANnpp environment server for state information, such as location and RGB camera images, fed this state input to the Q network, and subsequently sent a request to the iFANnpp server to move according to the maximum predicted expected discounted sum of reward.
The following reward function was used during training:
$\text{Reward}=\begin{cases}100,&\text{If Target Reached}\\ -100,&\text{If Collided}\\ 0,&\text{If Max Steps Reached}\\ -\frac{\sqrt{{(x_{i}-x_{t})}^{2}}}{x_{s}-x_{t}}&\text{Otherwise}\end{cases}$
Where $x_{i}$ is the current location ($x$, $y$, $z)$, $x_{t}$ is the target location, and $x_{s}$ is the start location.
The model achieved %100 success rate of reaching the target. When the radiation source was turned off, the agent moved through the radiation source and straight to the target. When the radiation source was turned on, the model navigated around the radiation source by moving around the source from the left side and reaching the target without touching the radiation source.
V-C1 Interpretability
In addition to observing the strategy that the robot learned, model interpretation methods were utilized to gain a sense for what visual features the robot learned. GradientShap was used, an importance attribution method, to qualitatively measure the salient features learned by the model [37].
Figure 10 displays both original images and GradientShap-overlayed images showing the most relevant pixels per image. The pixels are colored by importance magnitude, where blue corresponds to a higher importance value and white/gray corresponds to a lower importance value. In general, the model seems to be most affected by pixels surrounding relevant objects, possibly indicating that it is using edge information in order to detect these objects in the environment.
This demonstration shows that it is possible to use the iFANnpp platform in order to train a robotic task via reinforcement learning, demonstrating how all relevant features function together effectively.
Figure 10: Original and GradientShap Images of Relevant Objects for the Navigation Task
@@ VI Profiling
The main objective of the development of iFANnpp is to provide a platform for researchers to investigate various robotics research in nuclear power plants quickly and easily. It is essential to validate and demonstrate this tool’s effectiveness, efficiency, and reliability. To achieve this, a series of profiling tests were conducted to measure the run-time performance of various platform operations, along with memory and CPU usage.
@@
Baseline - 3.37 G 108.757% -
GPWR Communication(set) 18.160 ms 3.715 G 85.517% Passed (100%)
GPWR Communication(get) 2.37 ms 3.633 G 426.397% Passed (100%)
Unreal-Python Bridge (Read Image) 27.785 ms 4.557 G 235.518% Passed (100%)
Unreal-Python Bridge (Environment Step) 56.7 ms 4.590 G 255.583% Passed (100%)
Unreal-Python Bridge (Restart Episode) 104.977 ms 4.612 G 255.067% Passed (100%)
Unreal-Python Bridge (Move 2D) 20.928 ms 4.6 G 233.867% Passed (100%)
Thermal Vision 87.623 ms 4.424 G 186.062% Passed (100%)
Real-Time Top-Down Perspective 690.545 ms 3.662 G 76.868% Passed (100%)
Virtual Reality 720.47 ms 4.447 G 273.563 % Passed (100%)
@@ VI-A Speed Tests
Separate profiling was performed for the read and write operations as part of the GPWR communication test. The read operation queried the current state of the GPWR variables, while the write operation updated the state of GPWR variables. For both tests, 100 GPWR variables were transmitted between the iFANnpp client and the GPWR server. Each read and write operation was performed 100 times, with the execution time recorded in milliseconds. This test quantified the time required for the iFANnpp platform to synchronize the variables with GPWR. The average execution times for read and write operations are detailed in the table.
@@
The Python-Unreal bridge navigation commands test measured the runtime for the following independent operations: (1) capturing a first-person (robot perspective) image from iFANnpp via the Python client, (2) moving the robot in two dimensions (e.g., turning left or right by 15°, or moving forward or backward by 1 meter), with actions chosen randomly, and (3) restarting the Unreal Engine level or training episode. Additionally, the total time to complete a single ”step” operation was profiled, as defined by the Gymnasium v0.29.1 API [38]. The ”step” function involved moving the robot, capturing a first-person image, and calculating the reward. Each navigation command operation was executed 100 times, and the average execution time is detailed in the table.
@@
The thermal vision profile test measured the time required to visually update a component in the environment with thermal information, excluding the time taken to retrieve this information from GPWR. Thermal updates were manually triggered using a pre-programmed key press, and the total time for each update operation was recorded. The average time was calculated over 100 updates and is reported in the table.
@@
The real-time top-down test evaluated the time needed to capture a sequence of top-down images at fixed time intervals. A single operation was performed to capture 100 consecutive images, with a one-second delay between captures. The average time across 100 images is recorded as the final value in the table. The one-second delays were subtracted from the total capture times to isolate the operational overhead.
@@
The VR performance test measured the time required to execute a move command in a virtual reality environment, covering both the time taken to send the command from the client to the server and the time to receive the corresponding response. Each move command instructed the VR entity to navigate within the simulated environment, performing a basic action to move forward or backward by a specified distance. A total of 100 move commands were executed and the execution time for each command was recorded in milliseconds. The average execution time across all 100 commands is detailed in the table.

### v3

@@ IV-E Robot Navigation Utilities
Figure 4: Navigation Utility
iFANnpp includes functionality for visualizing the path that a robot takes. This feature is useful for robot debugging purposes, allowing researchers to visually compare a robot’s expected path against the path it took in practice (See Figure 4). Researchers can also identify potential patterns in the path that the robot is taking, improving the interpretability of its learned algorithm and directing the researcher toward potential algorithmic or mechanical modeling modification to improve the robot’s path planning and execution performance.
The path visualization feature is implemented in Unreal Engine as a component that can be added to a robot actor. A boolean variable can be selected through the editor to enable or disable the visualization feature. While enabled, the robot path is rendered onto the map. The robot path is implemented using a series of debug lines generated via calls to UE5 ”DrawDebugLine”. Successive points along the robot path are linearly interpolated to produce a smooth visual path that trails behind the robot. Both the thickness and color of the path can be customized through the editor.
Additionally, a boolean variable is exposed through the UE5 editor to allow for recording the robot location. The current robot location is recorded for every tick within the simulator and is logged to a CSV file along with the current timestamp in milliseconds. This logging feature can plot and visualize the robot path offline, independent of the iFANnpp environment, which can aid in post-hoc algorithm analysis. This feature can also be used to collect robot path demonstration trajectories to train robot agents, e.g., using behavioral cloning. Both human and machine policies can be used to produce demonstrations, as the robot can either be manually or pragmatically controlled.
To enhance the interpretability of the robot path in the first-person perspective, a compass feature is developed. This feature can also aid researchers in reviewing and debugging first-person robot navigation task experimentation. The compass features can be enabled or disabled interactively by pressing a keyboard key while the digital twin simulation is running. This displays an overhead icon that tracks and highlights the target relative to the robot, indicating the relative angle of the target (see Figure 4). This can inform a practitioner on how to potentially adjust their algorithm to improve performance on a navigation task.
@@ IV-F Real-Time Top-Down Perspective
A top-down camera class was created within Unreal Engine, inheriting from the camera actor class. This camera is positioned at a location in which the entire map can be captured. It becomes active through a single key input. This feature allows switching between first-person and top-down perspectives. The change in perspective is useful as it provides a comprehensive overview of the environment, shown in Figure 5. The top-down images are saved at predefined intervals to record critical moments and movements.
This feature offers several benefits. First, the top-down feature provides tracing the robot’s path. This perspective displays a clear visualization of the route taken by the robot, which helps researchers to understand the robot’s movement patterns. Second, this feature enables the precise identification of multiple robots that work simultaneously in the environment. In scenarios where coordination and interaction between several robots are critical, this feature can be utilized effectively. Third, the feature allows for comprehensive process monitoring of the entire nuclear power plant. Researchers can view the environment from above and track the interactions between various elements. Lastly, the feature captures a series of images throughout an operation at predefined intervals. The saving of images establishes a continuous record of the robot’s activities, which can be used for operation analysis, troubleshooting, and performance optimization.
@@ IV-G Thermal Vision
A thermal vision feature was implemented using ”Material” function from Unreal Engine. Thermal values are linearly interpolated between $100-$ 100 to 100100100 Celsius, visually represented by interpolating from blue to red. Actors without temperature properties are visually distinguished by being rendered in green, effectively removing them from the blue and red spectrum. This color exclusion process eliminates confusion with Unreal Engine actors that are naturally red or blue. As shown in Figure 6, this method provides a clear thermal view of the circulating water system, allowing users to interpret the temperature distribution easily. With real-time mirroring, this feature dynamically reflects the constantly changing environment and displays the entire map in thermal view.
This feature has many potential applications. First, the thermal view provides an intuitive and immediate understanding of temperature distributions across the plant. As mentioned, robots can make quicker decisions without the need to physically walk up to individual sensors. Second, this feature can enable comprehensive plant monitoring, ensuring that all critical components are consistently observed. Third, thermal images can be synthetically generated to address the lack of open-source visual thermal data in nuclear power plants. Moreover, the generated thermal data can uncover patterns, trends, and correlations that are difficult to observe in tabular datasets. While tabular data often require thorough analysis to identify subtle trends, this feature captures complex and minor variations in the environment, enabling the perception of multiple nuanced patterns simultaneously.
@@ V Use Cases
In this section, a representative set of realistic scenarios is provided, utilizing iFANnpp for future research projects. These scenarios demonstrate multiple research directions that could take advantage of the features provided by the iFANnpp platform.
@@ V-A Example Scenario: Swarm of Robots
One valuable avenue for robotics research in nuclear power plants (as well as industrial systems in general) is swarm robotics [34, 35]. Robotics swarms allow for the orchestration of a relatively large number of simplistic robots capable of coordinating together to perform complex behavior. Robotic swarm deployment can be used to form flexible functional configurations, create mobile assembly lines, and perform coordinated maintenance and repair tasks.
To demonstrate the viability of simulating a robotic swarm, a swarm of twenty TurtleBots was placed within the turbine hall of the iFANnpp environment. The robots were placed adjacent to one another in a closely clustered configuration. Figure 7 displays the swarm robot simulation. The left-most and right-most images display first-person camera views from two individual robots within the swarm. Each robot was individually controlled from these views. The middle image is an overhead display of the twenty-bot swarm. By setting net mode within the Unreal Editor to ”Play as Listen Server” and extending the number of players to 2, multiple turtle-bots were controlled using the mouse and keyboard. Control was transferred from one robot to another. As with other simulation tasks, iFANnpp allows for iterating on both the mechanical and algorithmic components of swarm robot systems.
@@
V-B Example Scenario: Robot Remote Control Via Digital Twin
The implementation of VR technology within the digital twin, especially in critical infrastructure like nuclear power plants, provides significant value. Teleoperation, remotely control of robots using the VR technology, is a key method for ensuring safe operation. In the scenario illustrated in Figure 8, a teleoperation setup is demonstrated in the turbine hall. The VR allows an operator to control a robot remotely and navigate the robot to locate and close a valve to prevent the system from overflooding.
This application demonstrates how VR can increase safety, enabling operators to address emergencies without being physically present in dangerous areas. Furthermore, VR improves operator remote working accuracy and reduced cognitive load through enhanced situational awareness.
There are multiple research directions that can utilize the integration of VR technology with nuclear power plants. Scenarios such as mobile robot teleoperation for pipe maintenance [36] and training platform development [37] have been experimentally validated, and the corresponding results are reported in the references cited. Other topics, including advanced human-robot collaboration and disaster preparedness simulations, remain conceptual at this stage and are subjects of ongoing and planned research.
@@ V-C Example Scenario: Robot Navigation
Research has focused extensively on enhancing the use of robots in industrial systems, including work focusing on their movement and decision-making capabilities [38]. One critical feature is their ability to navigate effectively, which involves selecting optimal paths while avoiding hazardous areas, such as those with high radiation levels in nuclear power plants. While robot teleportation allows for manual human control, autonomous robot navigation can be valuable when a human is not present or able to operate a robot remotely (e.g., if it is difficult to establish wireless communication). In addition, autonomous navigation capabilities can remove the burden on human operators for mundane mobile tasks such as fetching a tool that the operator needs. For example, robots are utilized in autonomous rapid radiation reconstruction in the work [39], which could be tested in iFANnpp. Therefore, research investigating the design and improvement of autonomous robot navigation algorithms is of interest.
Integration of the Python-Unreal bridge with iFANnpp enables the development of novel robot navigation algorithms, critical for mobile robot deployment within an NPP. State information, such as location and RGB camera images, can be queried from the environment and used to inform a robot policy to perform intelligent navigation, including reinforcement-learning-driven control. With this workflow, many scenarios can be evaluated, such as physically constrained navigation with unanticipated environmental changes.
V-D Example Scenario: Cyber-Physical Security Testing
With the rapid digitalization and emerging technology adoption in nuclear field, cybersecurity requires more attention and further research [40]. Previous works have employed the use of the GPWR simulator for cybersecurity testing. Yoo et al. [41] utilized GPWR to simulate a false-data-injection attack in order to design an automated attack mitigation strategy using autoencoder output to correct for the attack effects. Chen et al. [42] used GPWR as the backbone for hardware-in-the-loop testing and conducted a supply chain attack simulation for future cyber-attack development and NPP system robustness testing. While the data from this GPWR can be used to test different cyber-attack scenarios, integration of the simulator with the physically modeled iFANnpp plant allows for additionally incorporating physical information with security-focused experiments. This enables the development of multimodal cyber-attack prevention or detection techniques - both physical information (such as from images), as well as numerical plant sensor values can be used to identify potential threats to the plant to prevent damage.
Example scenarios can vary by testing different degrees of physical and cyber-related consequences and effects. Physical tracking can be used to identify if an unauthorized entity is located within the plant. Correlation of process data with observed robot or human function can help to identify when physical interaction with the plant is benign or malicious.
While humans present one physical attack vector, it is also critical to ensure the security of the robots themselves for practical real-world robot deployment for monitoring and maintenance tasks within an NPP. Additional tests can be conducted such as identifying if a robot is being maliciously controlled by combining technologies such as action identification and anomaly detection with plant process data. Unintentional robot malfunction identification technique development can also be conducted.
@@ VI Profiling
The main objective of the development of iFANnpp is to provide a platform for researchers to investigate various robotics research in nuclear power plants quickly and easily. Validating and demonstrating the effectiveness, efficiency, and reliability of the platform is essential. To achieve this, a series of profiling tests were conducted to measure the runtime performance of various platform operations, along with memory and CPU usage.
@@
Baseline - 3.37 GB 108.76% -
GPWR Communication(set) 18.16 ms 3.72 GB 85.52% Passed (100%)
GPWR Communication(get) 2.37 ms 3.63 GB 426.40% Passed (100%)
Unreal-Python Bridge (Read Image) 27.79 ms 4.56 GB 235.52% Passed (100%)
Unreal-Python Bridge (Environment Step) 56.70 ms 4.59 GB 255.58% Passed (100%)
Unreal-Python Bridge (Restart Episode) 104.98 ms 4.61 GB 255.07% Passed (100%)
Unreal-Python Bridge (Move 2D) 20.93 ms 4.60 GB 233.87% Passed (100%)
Thermal Vision 87.62 ms 4.42 GB 186.06% Passed (100%)
Real-Time Top-Down Perspective 690.55 ms 3.66 GB 76.87% Passed (100%)
Virtual Reality 720.47 ms 4.45 GB 273.56 % Passed (100%)
@@ VI-A Speed Tests
Separate profiling was performed for the read and write operations as part of the GPWR communication test. The read operation queried the current state of the GPWR variables, while the write operation updated the state of the GPWR variables. For both tests, 100 GPWR variables were transmitted between the iFANnpp client and the GPWR server. Each read and write operation was performed 100 times, with the execution time recorded in milliseconds. This test quantified the time required for the iFANnpp platform to synchronize the variables with GPWR. The average execution times for read and write operations are detailed in the table. The write operation yielded an average runtime of 18.16 milliseconds the read operation averaged 2.37 milliseconds.
@@
The Python-Unreal bridge navigation commands test measured the runtime for the following independent operations: (1) capturing a first-person (robot perspective) image from iFANnpp via the Python client, (2) moving the robot in two dimensions (e.g., turning left or right by 15°, or moving forward or backward by 1 meter), with actions chosen randomly, and (3) restarting the Unreal Engine level or training episode. Additionally, the total time to complete a single ”step” operation was profiled, as defined by the Gymnasium v0.29.1 API [43]. The ”step” function involved moving the robot, capturing a first-person image, and calculating the reward. Each navigation command operation was executed 100 times, and the average execution time is detailed in the table. These operations were profiled as they can be used as fundamental operations for the purposes of reinforcement-learning-based navigation. The most expensive operation was restarting the episode, which took 104.977 milliseconds on average, while reading a first-person image and moving one step averaged 27.79 and 20.93 milliseconds, respectively. The total time cost per environment step was 56.7 milliseconds.
@@
The thermal vision profile test measured the time required to visually update a component of the environment with thermal information, excluding the time taken to retrieve this information from GPWR. Thermal updates were manually triggered using a pre-programmed key press, and the total time for each update operation was recorded. The average time was calculated over 100 updates and is reported in the table, which took an average of 87.62 milliseconds per operation.
@@
The real-time top-down test evaluated the time needed to capture a sequence of top-down images at fixed time intervals. A single operation was performed to capture 100 consecutive images, with a one-second delay between captures. The average time across 100 images is recorded as the final value in the table. The one-second delays were subtracted from the total capture times to isolate the operational overhead. The average time per 100 recorded images was 690.55 milliseconds. Note that this operation can be performed concurrently with other operations such as robot navigation.
@@
The VR performance test measured the time required to execute a move command in a VR environment, covering both the time taken to send the command from the client to the server and the time to receive the corresponding response. Each move command instructed the VR entity to navigate within the simulated environment, performing a basic action to move forward or backward by a specified distance. A total of 100 move commands were executed and the execution time for each command was recorded in milliseconds. The average execution time across all 100 commands yielded a runtime of 720.47 millisecond average.
The results of the speed test demonstrate that the digital twin is applicable to multiple research directions. The relatively low latency of the platform indicates that several features can be used without critical delays. This is particularly important for AI model training, which may require substantial computational resources on top of the iFANnpp overhead. The results therefore signify that the platform is suitable for model training as well.